# Leveraging the Bias-Variance Tradeoff in Quantum Chemistry for Accurate Negative Singlet-Triplet Gap Predictions: A Case for Double-Hybrid DFT

### v2

@@ 3 Bias Correction for Low-Variance Predictors
    Let $\hat{y}_{1}$ and $\hat{y}_{2}$ denote two predictive models (or predictors) for a given property. In this study, they represent two parameterizations of a DFA used for determining STG. Further, $\hat{y}_{\rm true}$ represents the exact theoretical value of the property under consideration. The best available finite-data estimate of $\hat{y}_{\rm true}$ is denoted as the reference value, $\hat{y}^{\rm ref}$. The $k$-th prediction from $\hat{y}_{1}$ and the corresponding reference value are denoted as $\hat{y}_{1,k}$ and $\hat{y}_{{\rm ref},k}$, respectively.
@@
    The expected value (i.e., mean) of $\hat{y}_{1}$ is given by $\mathbb{E}[\hat{y}_{1}]$. The variance of $\hat{y}_{1}$ around its mean is defined as: ${\rm Var}(\hat{y}_{1})=\mathbb{E}[\hat{y}_{1}^{2}]-\mathbb{E}[\hat{y}_{1}]^{2}$. The covariance between two predictors measures their linear relationship and is defined as: ${\rm Cov}(\hat{y}_{1},\hat{y}_{2})=\mathbb{E}[\hat{y}_{1}\hat{y}_{2}]-\mathbb{E}[\hat{y}_{1}]\mathbb{E}[\hat{y}_{2}]$.
@@
    The error metrics are defined as follows. Mean signed error (MSE): ${\rm MSE}(\hat{y}_{1})=\mathbb{E}[\hat{y}_{1}-\hat{y}_{\rm ref}]$, mean absolute deviation (MAD): ${\rm MAD}(\hat{y}_{1})=\mathbb{E}[|\hat{y}_{1}-\hat{y}_{\rm ref}|]$, standard deviation of error (SDE) for a sample of size $N$: ${\rm SDE}(\hat{y}_{1})=\sqrt{\frac{1}{N-1}\sum_{k=1}^{N}(\hat{y}_{1,k}-\hat{y}_{{\rm ref},k})^{2}}$. In the limit of a large sample (i.e., for a population), the square of SDE corresponds to the variance.
@@
    The term ‘bias’ refers to a systematic shift in a distribution and requires a reference for comparison. Throughout this study, bias specifically refers to systematic shifts in prediction error: $\hat{y}_{1}-\hat{y}_{\rm ref}$. Similarly, we distinguish between: ${\rm Var}(\hat{y}_{1})$ and ${\rm Var}(\hat{y}_{1}-\hat{y}_{\rm ref})$, where the former indicates spread around the mean value of $\hat{y}_{1}$ while latter quantifies the spread of its prediction errors, which is the primary focus of this study.
Bias Correction via Ordinary Least Squares Regression:
The bias correction scheme applies in cases where two predictors, $\hat{y}_{1}$ and $\hat{y}_{2}$, are available with known bias and variance relative to the reference values, $\hat{y}^{\rm ref}$. We assume that $\hat{y}_{1}$ has low variance in prediction errors (i.e., small SDE) but high bias (i.e., large magnitude of MSE), and $\hat{y}_{2}$ has low bias but high variance in prediction errors. These conditions are expressed as:
$|\mathbb{E}\left[\hat{y}_{1}-\hat{y}^{\rm ref}\right]|>0,\,\text{Var}\left(\hat{y}_{1}-\hat{y}^{\rm ref}\right)\approx 0;\qquad\mathbb{E}\left[\hat{y}_{2}-\hat{y}^{\rm ref}\right]\approx 0,\,\text{Var}\left(\hat{y}_{2}-\hat{y}^{\rm ref}\right)>0.$ (3.1)
We define a new estimator $\hat{y}_{1}^{*}$ by linearly scaling $\hat{y}_{1}$:
$\hat{y}_{1}^{*}$ $=$ $a\hat{y}_{1}+b,$ (3.2)
where $a$ and $b$ are determined via ordinary least squares regression, minimizing the squared error with respect to $\hat{y}_{2}$:
$\min_{a,b}\sum_{i}\left(a\hat{y}_{1,i}+b-\hat{y}_{2,i}\right)^{2}.$ (3.3)
The optimal values of $a$ and $b$ are given by:
$a=\frac{\text{Cov}(\hat{y}_{1},\hat{y}_{2})}{\text{Var}(\hat{y}_{1})};\quad b=\mathbb{E}[\hat{y}_{2}]-a\mathbb{E}[\hat{y}_{1}].$ (3.4)
Error analysis:
We want to determine whether the bias and variance of $\hat{y}_{1}^{*}$ match those of $\hat{y}_{2}$ and $\hat{y}_{1}$, respectively. The mean of $\hat{y}_{1}^{*}$ is:
$\mathbb{E}[\hat{y}_{1}^{*}]=\mathbb{E}[a\hat{y}_{1}+b]=a\mathbb{E}[\hat{y}_{1}]+b.$ (3.5)
Substituting the expression for $b$ from Eq. 3.4, we find:
$\mathbb{E}[\hat{y}_{1}^{*}]=a\mathbb{E}[\hat{y}_{1}]+\mathbb{E}[\hat{y}_{2}]-a\mathbb{E}[\hat{y}_{1}]=\mathbb{E}[\hat{y}_{2}].$ (3.6)
Thus, least-squares regression guarantees that the mean of $\hat{y}_{1}^{*}$ aligns with that of the reference—low-bias predictor $\hat{y}_{2}$—ensuring that the bias is corrected. The MSE of $\hat{y}_{1}^{*}$ follows:
$\mathbb{E}[\hat{y}_{1}^{*}-\hat{y}^{\rm ref}]$ $=$ $\mathbb{E}[a\hat{y}_{1}+b-\hat{y}^{\rm ref}]=\mathbb{E}[a\hat{y}_{1}]+\mathbb{E}[b]-\mathbb{E}[\hat{y}^{\rm ref}]$ (3.7)
@@
The variance of $\hat{y}_{1}^{*}$ follows:
$\text{Var}(\hat{y}_{1}^{*})=\text{Var}(a\hat{y}_{1}+b)=a^{2}\text{Var}(\hat{y}_{1}).$ (3.8)
@@
${\rm Var}\left(\hat{y}_{1}^{*}-\hat{y}^{\rm ref}\right)$ $=$ ${\rm Var}\left(a\hat{y}_{1}+b-\hat{y}^{\rm ref}\right).$ (3.9)
Applying the variance sum formula, ${\rm Var}\left(A+B\right)={\rm Var}\left(A\right)+{\rm Var}\left(B\right)+{\rm Cov}\left(A,B\right)$, we arrive at:
${\rm Var}\left(\hat{y}_{1}^{*}-\hat{y}^{\rm ref}\right)$ $=$ ${\rm Var}\left(a\hat{y}_{1}\right)+{\rm Var}\left(b-\hat{y}^{\rm ref}\right)-2{\rm Cov}\left(a\hat{y}_{1},b-\hat{y}^{\rm ref}\right)$ (3.10)
$=$ ${\rm Var}\left(a\hat{y}_{1}\right)+{\rm Var}\left(\hat{y}^{\rm ref}\right)-2{\rm Cov}\left(a\hat{y}_{1},\hat{y}^{\rm ref}\right)$
$=$ ${\rm Var}\left(a\hat{y}_{1}-\hat{y}^{\rm ref}\right)$
$=$ $a^{2}{\rm Var}\left(\hat{y}_{1}-\hat{y}^{\rm ref}/a\right).$
@@
Practical Considerations:
@@
4 Results and Discussions

### v3

@@ III.1 Definitions
    Let $\hat{y}_{1}$ and $\hat{y}_{2}$ denote two predictive models (or predictors) for a given property. In this study, they represent two parameterizations of a dh-DFA used for determining STG. Further, $\hat{y}^{\rm true}$ represents the exact theoretical value of the property under consideration. The best available finite-data estimate of $\hat{y}^{\rm true}$ is denoted as the reference value, $\hat{y}^{\rm ref}$. The $k$-th prediction from $\hat{y}_{1}$ and the corresponding reference value are denoted as $\hat{y}_{1,k}$ and $\hat{y}_{{\rm ref},k}$, respectively.
@@
    The expected value (i.e., mean) of $\hat{y}_{1}$ is given by $\mathbb{E}[\hat{y}_{1}]$. The variance of $\hat{y}_{1}$ around its mean is defined as ${\rm Var}(\hat{y}_{1})=\mathbb{E}[\hat{y}_{1}^{2}]-\mathbb{E}[\hat{y}_{1}]^{2}$. The covariance between two predictors measures their linear relationship and is defined as ${\rm Cov}(\hat{y}_{1},\hat{y}_{2})=\mathbb{E}[\hat{y}_{1}\hat{y}_{2}]-\mathbb{E}[\hat{y}_{1}]\mathbb{E}[\hat{y}_{2}]$.
@@
    The error metrics are defined as follows. Mean signed error (MSE) is defined as ${\rm MSE}(\hat{y}_{1})=\mathbb{E}[\hat{y}_{1}-\hat{y}_{\rm ref}]$, while mean absolute deviation (MAD) is defined as ${\rm MAD}(\hat{y}_{1})=\mathbb{E}[|\hat{y}_{1}-\hat{y}_{\rm ref}|]$ Standard deviation of error (SDE) for a sample of size $N$ is defined as ${\rm SDE}(\hat{y}_{1})=\sqrt{\frac{1}{N-1}\sum_{k=1}^{N}(\hat{y}_{1,k}-\hat{y}_{{\rm ref},k}-{\rm MSE})^{2}}$. In the limit of a large sample (i.e., for a population), the square of SDE corresponds to the variance of the error ${\rm Var}(\hat{y}_{1}-\hat{y}_{\rm ref})$.
@@
    The term ‘bias’ refers to a systematic shift in a distribution and requires a reference for comparison. Throughout this study, bias specifically refers to systematic shifts in prediction error $\hat{y}_{1}-\hat{y}_{\rm ref}$. Similarly, we distinguish between ${\rm Var}(\hat{y}_{1})$ and ${\rm Var}(\hat{y}_{1}-\hat{y}_{\rm ref})$, where the former indicates spread around the mean value of $\hat{y}_{1}$ but the latter quantifies the spread of its prediction errors, which is the primary focus of this study.
III.2 Bias Correction via Ordinary Least Squares Regression
The bias correction scheme is applicable in cases where two predictors, $\hat{y}_{1}$ and $\hat{y}_{2}$, are available with known bias and variance determined a priori using reference values, $\hat{y}^{\rm ref}$. For example, the bias and variance of the two models can be determined using a subset of the entire dataset.
We assume that $\hat{y}_{1}$ has low variance in prediction errors (i.e., small SDE) but high bias (i.e., large magnitude of MSE), and $\hat{y}_{2}$ has low bias but high variance in prediction errors. These conditions are expressed as
$|\mathbb{E}\left[\hat{y}_{1}-\hat{y}^{\rm ref}\right]|$ $>$ $0,\,\text{Var}\left(\hat{y}_{1}-\hat{y}^{\rm ref}\right)\approx 0\,\text{and}$
$\mathbb{E}\left[\hat{y}_{2}-\hat{y}^{\rm ref}\right]$$\approx$ $0\displaystyle 0,\,\text{Var}\left(\hat{y}_{2}-\hat{y}^{\rm ref}\right)>$ 0 . (1)
We define a new estimator $\hat{y}_{1}^{*}$ by linearly scaling $\hat{y}_{1}$
$\hat{y}_{1}^{*}$ $=$ $a\hat{y}_{1}+b,$ (2)
where $a$ and $b$ are determined via ordinary least squares regression, minimizing the squared error
$\min_{a,b}\sum_{i}\left(a\hat{y}_{1,i}+b-\hat{y}_{2,i}\right)^{2}$ (3)
with respect to $\hat{y}_{2}$, which is used as an internal reference. The optimal values of $a$ and $b$ are given by
$a=\frac{\text{Cov}(\hat{y}_{1},\hat{y}_{2})}{\text{Var}(\hat{y}_{1})};\quad b=\mathbb{E}[\hat{y}_{2}]-a\mathbb{E}[\hat{y}_{1}].$ (4)
III.3 Error analysis
We want to determine whether the bias and variance of $\hat{y}_{1}^{*}$ match those of $\hat{y}_{2}$ and $\hat{y}_{1}$, respectively. The mean of the corrected model, $\hat{y}_{1}^{*}$, is
$\mathbb{E}[\hat{y}_{1}^{*}]=\mathbb{E}[a\hat{y}_{1}+b]=a\mathbb{E}[\hat{y}_{1}]+b.$ (5)
Substituting the expression for $b$ from Eq. 4, we find:
$\mathbb{E}[\hat{y}_{1}^{*}]=a\mathbb{E}[\hat{y}_{1}]+\mathbb{E}[\hat{y}_{2}]-a\mathbb{E}[\hat{y}_{1}]=\mathbb{E}[\hat{y}_{2}].$ (6)
Thus, least-squares regression ensures that the mean of $\hat{y}_{1}^{*}$ aligns with that of the reference, the low-bias predictor $\hat{y}_{2}$, thereby correcting the bias.
The MSE of $\hat{y}_{1}^{*}$ is
$\mathbb{E}[\hat{y}_{1}^{*}-\hat{y}^{\rm ref}]$ $=$ $\mathbb{E}[a\hat{y}_{1}+b-\hat{y}^{\rm ref}]=\mathbb{E}[a\hat{y}_{1}]+\mathbb{E}[b]-\mathbb{E}[\hat{y}^{\rm ref}]$ (7)
@@ III.3 Error analysis
The variance of $\hat{y}_{1}^{*}$ can be derived as
$\text{Var}(\hat{y}_{1}^{*})=\text{Var}(a\hat{y}_{1}+b)=a^{2}\text{Var}(\hat{y}_{1}).$ (8)
@@
${\rm Var}\left(\hat{y}_{1}^{*}-\hat{y}^{\rm ref}\right)$ $=$ ${\rm Var}\left(a\hat{y}_{1}+b-\hat{y}^{\rm ref}\right).$ (9)
Applying the variance sum formula, ${\rm Var}\left(A+B\right)={\rm Var}\left(A\right)+{\rm Var}\left(B\right)+{\rm Cov}\left(A,B\right)$, we arrive at
${\rm Var}\left(\hat{y}_{1}^{*}-\hat{y}^{\rm ref}\right)$ $=$ $a^{2}{\rm Var}\left(\hat{y}_{1}-\hat{y}^{\rm ref}/a\right).$ (10)
@@
Table 1: Excitation energies of S1 and T1 levels of the twelve benchmark systems FIG 1, along with the corresponding STGs, are presented at different theoretical levels using the aug-cc-pVTZ basis set. For reference, TBE results are also provided, which are based on composite estimations. Error metrics and the reference TBE values are also provided. MSE: mean signed error, MAD: mean absolute deviation, and SDE: standard deviation of the error. Wherever necessary, the convention for deviation is ‘Method--TBE.’ All values are in eV.
# L-CC2a CC2b L-ADC(2)a ADC(2)b TBEb
S1 T1 S1-T1 S1 T1 S1-T1 S1 T1 S1-T1 S1 T1 S1-T1 S1 T1 S1-T1
1 $2.759$ $2.995$ $-0.236$ $2.767$ $3.006$ $-0.239$ $2.680$ $2.934$ $-0.254$ $2.675$ $2.921$ $-0.246$ $2.717$ $2.936$ $-0.219$
2 $1.043$ $1.174$ $-0.131$ $1.051$ $1.181$ $-0.130$ $1.001$ $1.138$ $-0.137$ $1.001$ $1.138$ $-0.137$ $0.979$ $1.110$ $-0.131$
3 $1.606$ $1.713$ $-0.107$ $1.615$ $1.721$ $-0.106$ $1.553$ $1.666$ $-0.113$ $1.551$ $1.664$ $-0.113$ $1.562$ $1.663$ $-0.101$
4 $2.220$ $2.354$ $-0.134$ $2.235$ $2.366$ $-0.131$ $2.167$ $2.300$ $-0.133$ $2.159$ $2.298$ $-0.139$ $2.177$ $2.296$ $-0.119$
5 $2.162$ $2.286$ $-0.124$ $2.178$ $2.296$ $-0.118$ $2.097$ $2.228$ $-0.131$ $2.098$ $2.225$ $-0.127$ $2.127$ $2.230$ $-0.103$
6 $0.896$ $0.985$ $-0.089$ $0.903$ $0.988$ $-0.085$ $0.852$ $0.951$ $-0.099$ $0.851$ $0.945$ $-0.094$ $0.833$ $0.904$ $-0.071$
7 $0.755$ $0.825$ $-0.070$ $0.762$ $0.827$ $-0.065$ $0.709$ $0.787$ $-0.078$ $0.708$ $0.782$ $-0.074$ $0.693$ $0.735$ $-0.042$
8 $0.612$ $0.665$ $-0.053$ $0.623$ $0.681$ $-0.058$ $0.565$ $0.630$ $-0.065$ $0.565$ $0.635$ $-0.070$ $0.554$ $0.583$ $-0.029$
9 $1.321$ $1.540$ $-0.219$ $1.343$ $1.555$ $-0.212$ $1.267$ $1.488$ $-0.221$ $1.274$ $1.488$ $-0.214$ $1.264$ $1.463$ $-0.199$
10 $1.578$ $1.907$ $-0.329$ $1.597$ $1.916$ $-0.319$ $1.525$ $1.846$ $-0.321$ $1.529$ $1.840$ $-0.311$ $1.522$ $1.827$ $-0.305$
11 $0.828$ $1.025$ $-0.197$ $0.866$ $1.054$ $-0.188$ $0.786$ $0.997$ $-0.211$ $0.799$ $0.990$ $-0.191$ $0.779$ $0.974$ $-0.195$
12 $1.246$ $1.179$ $+0.067$ $1.275$ $1.207$ $+0.068$ $1.248$ $1.177$ $+0.071$ $1.262$ $1.179$ $+0.083$ $1.209$ $1.147$ $+0.062$
MSE $+0.051$ $+0.065$ $-0.014$ $+0.067$ $+0.077$ $-0.011$ $+0.003$ $+0.023$ $-0.020$ $+0.005$ $+0.020$ $-0.015$
MAD $+0.051$ $+0.065$ $+0.015$ $+0.067$ $+0.077$ $+0.013$ $+0.017$ $+0.023$ $+0.022$ $+0.021$ $+0.023$ $+0.019$
SDE $+0.010$ $+0.016$ $+0.014$ $+0.011$ $+0.013$ $+0.011$ $+0.021$ $+0.018$ $+0.013$ $+0.025$ $+0.020$ $+0.016$
  • a This work
  • b From Ref. Loos, Lipparini, and Jacquemin, 2023. Corrected CC2 and ADC(2)values for System 10 are from Ref. Loos, Lipparini, and Jacquemin, 2025.
III.4 Practical Considerations
@@
The internally referenced bias correction procedure discussed above is general and can be applied to various scenarios. The main ingredient to using this procedure to solve various quantum chemistry problems lies in obtaining the bias and variance of two formalisms of a method (such as a DFA) through benchmarking against reference data. A more critical aspect is the availability of a suitable $\hat{y}_{1}$ and $\hat{y}_{2}$ predictors to test the hypothesis presented here.
It is often the case that for molecular properties, $\hat{y}_{2}$ predictors with low bias (but with high variance) are more commonly reported. For example, it is well known that semi-empirical methods, such as PM6 or PM7, are fitted to reference data to reduce bias. Hence, they often exhibit lower systematic errors compared to even the most popular DFAs. For example, for the curated dataset, PPE1694, comprising heats of formation of 1694 molecules, PM6 and PM7 achieved mean absolute deviations (MADs) of 4.02 and 3.89 kcal/mol Das, Chakraborty, and Ramakrishnan (2021). For the same set, B3LYP-hybrid-GGA achieved a higher MAD of 4.67 kcal/mol. However, predictors with low variance and high bias, $\hat{y}_{1}$, are seldom reported, partly because MAD is often preferred as a standard error metric over SDE. Hence, in most studies, the critical information regarding the variance that is better expressed in the form of SDE is missing. Further investigation is necessary to fully establish the generality of this approach and determine how it can be transferred and applied to a broader range of molecular systems.
IV Results and Discussions
IV.1 Performance of L-CC2 and L-ADC(2)
We begin our discussions with the analysis of the performance of L-CC2 and L-ADC(2), which have not been previously applied to model negative STGs. Table 1 presents the excitation energies and STGs of the twelve benchmark systems (shown in Figure 1) at the CC2 and ADC(2) method from Ref. Loos, Lipparini, and Jacquemin, 2023, along with results obtained with their local-correlated variants. All values are reported for the aug-cc-pVTZ basis set, which is considered to yield nearly converged S1 and T1 energies for small organic molecules such as the benchmark systems considered here. For comparison, we have also collected the TBE values from Ref. Loos, Lipparini, and Jacquemin, 2023, which we use as reference values to evaluate the accuracies of various methods.
As stated in Ref. Loos, Lipparini, and Jacquemin, 2023, structure 11 in Figure 1 (with $D_{3\rm h}$ symmetry) corresponds to a saddle point on the potential energy surface (PES) at the MP2/6-311G($d$,$p$) level, whereas the actual minimum energy geometry exhibits $C_{3\rm h}$ symmetry (structure 12 in Figure 1). Another study Majumdar et al. (2024) found that the high-symmetry forms of structures 6, 7, and 8 are not true minima, while structure 2 is a very shallow potential well at the CCSD(T)/cc-pVTZ level. Overall, four out of twelve structures of DFIST candidate molecules considered in this study are transition states; however, we include them in our benchmarking of STGs due to the availability of TBE-level results and the lack of alternative high-quality data.
Table 2: L-CC2 energies and L-ADC(2) energies with aug-cc-pVDZ of the S1 and T1 states with respect to the S0 ground state along with the singlet-triplet gap, S1-T1, of 12 triangular benchmark systems. All values are in eV and # indicates compound number as in Figure 1.
# L-CC2 L-ADC(2)
S1 T1 S1-T1 S1 T1 S1-T1
1 $2.745$ $2.992$ $-0.247$ $2.661$ $2.916$ $-0.255$
2 $1.050$ $1.178$ $-0.128$ $1.020$ $1.154$ $-0.134$
3 $1.602$ $1.711$ $-0.109$ $1.552$ $1.672$ $-0.12$
4 $2.211$ $2.341$ $-0.130$ $2.164$ $2.296$ $-0.132$
5 $2.147$ $2.274$ $-0.127$ $2.083$ $2.221$ $-0.138$
6 $0.902$ $0.988$ $-0.086$ $0.874$ $0.976$ $-0.102$
7 $0.757$ $0.826$ $-0.069$ $0.733$ $0.813$ $-0.080$
8 $0.613$ $0.665$ $-0.052$ $0.589$ $0.652$ $-0.063$
9 $1.304$ $1.519$ $-0.215$ $1.252$ $1.472$ $-0.220$
10 $1.563$ $1.888$ $-0.325$ $1.507$ $1.825$ $-0.318$
11 $0.834$ $1.021$ $-0.187$ $0.788$ $0.987$ $-0.199$
12 $1.257$ $1.185$ $+0.072$ $1.265$ $1.180$ $+0.085$
MSE +0.047 +0.060 --0.013 +0.006 +0.025 --0.019
MAD +0.047 +0.060 +0.016 +0.031 +0.030 +0.023
SDE +0.016 +0.017 +0.013 +0.035 +0.032 +0.017
Note that for system 10 (in Figure 1), STGs at ADC(2) and CC2 levels were originally reported as $-0.435$ eV and $-0.446$ eV, respectively Loos, Lipparini, and Jacquemin (2023). These values were later corrected Loos, Lipparini, and Jacquemin (2025) to $-0.311$ eV and $-0.319$ eV in better agreement with the TBE value of $-0.305$ eV. We considered these latter values in Table 1, amounting to MSE||||MAD||||SDE error metrics as $-0.011|0.013|0.011$ eV and $-0.015|0.019|0.016$ eV for STGs predicted with CC2 and ADC(2) compared to TBE. Even though the errors in STGs are lower for CC2 compared to ADC(2), as seen from Table 1, this is due to a favorable cancellation of errors in CC2. For the individual excitation energies, ADC(2) delivers lower error metrics compared to CC2.
While using the local-correlated variants, the errors for L-CC2 and L-ADC(2) were determined as $-0.014|0.015|0.014$ and $-0.020|0.022|0.013$ eV, respectively. These values are similar to those of the canonical methods. For both CC2 and ADC(2), introducing local correlation slightly increases the errors. However, the final errors are still sufficiently low to enable reliable determination of negative STGs with both L-CC2 and L-ADC(2). While we did not perform a speed test to determine the speedups due to local-correlation approximation, we observed L-CC2 and L-ADC(2) when using the smaller aug-cc-pVDZ basis set to enable excited-state calculations of molecules larger than the benchmark systems (such as those discussed in Section IV.5), which failed to converge successfully when using RI-based ADC(2) due to CPU memory constraints.
To explore the suitability of L-CC2 and L-ADC(2) for high-throughput data generation for identifying DFIST candidates, we performed calculations with the smaller aug-cc-pVDZ basis set. The corresponding results are collected in Table 2. L-CC2 and L-ADC(2) error metrics MSE||||MAD||||SDE determined with the aug-cc-pVDZ basis set $-0.013|0.016|0.013$ and $-0.019|0.023|0.017$ eV, respectively, are almost identical to the afore-stated values based on the larger aug-cc-pVTZ basis set ($-0.014|0.015|0.014$ and $-0.020|0.022|0.013$ eV, respectively). This ensures that the aug-cc-pVDZ basis set is sufficiently converged for further explorations.